A local animal feed company makes its feed by the ton, which is 2000 pounds. They want to include a medication in the feed. Each cow needs 300 milligrams (mg) of this medication a day and each cow consumes 15 pounds of feed per day. If there are 1,000 milligrams in a gram, how many grams of the medication should the feed company add for each ton of feed they produce.
Answer by richwmiller(17219)

300/3=100
15/3=5
100 milligrams for every 5 pounds
2000/5=400 doses
400*100=40,000 milligrams= 40 grams